# On Speeding Up Language Model Evaluation

### v2

@@ 4.4 Main Results
In Figure 3, we plot the performance of baselines and our algorithms on all six datasets (columns) as the budget $T$ increases from 5% to 100% of the total number of method-example pairs. In each row, we evaluate the algorithms on a different metric (either top 1 precision with different $\epsilon$, $p$ or NDCG@10). Each line in the figure is the average result over 50 independent trials with different seeds. For example, an average top 1 precision of 0.9 indicates that 45 out of 50 trials predict the best method correctly at the budget level that its x-coordinate represents. In addition, we calculate $H_{1}$ as defined in Corollary 1 that quantifies the difficulty of finding the best method on a dataset. Intuitively, a higher $H_{1}$ value suggests that the method set $\mathcal{F}$ is large or there are many methods that have similar performance with the best one and distinguishing them can be challenging. In Appendix A Figure 5, we also plot the histogram to show the distribution of performance among $\mathcal{F}$ on these datasets.
@@
As seen from Figure 3, both UCB-E and UCB-E-LRF consistently achieve high precision and NDCG with much less budget compared to the baselines. For example, on AlpacaEval (Drop Annotator), our proposed algorithms can reach precision of 1 with just 8% budget whereas the baselines require 80-90%, an order of magnitude more budget needed. These results suggest that it is entirely possible to identify the best method without exhaustively evaluating all method-example pairs. They further demonstrate that the active selection algorithms (our two algorithms) are more efficient than the non-active algorithms (three baselines). Additionally, the better NDCG performance from our proposed algorithms shows that our methods can more correctly rank top-performance methods.
@@ 4.5 More Empirical Analysis
Yes, going through Figure 3 from left to right, as $H_{1}$ decreases, the percentage of matrix evaluation needed to reach precision of 1 also generally decreases from more than 20% to just under 5%. Moreover, the $H_{1}$ values seem to be related to the tasks. That is, prompt engineering datasets typically have higher $H_{1}$ possibly due to the homogeneity of prompt performance with the same LLM. In contrast, datasets that benchmark different LLMs such as AlpacaEval is much easier to find the best performing model.
@@
To study the effect of selecting both the next method and the next example to evaluate using the uncertainty matrix $R$, we consider an ablation of UCB-E-LRF. In this variant, called UCB-E-LRF (Score Only), we modify the UCB-E algorithm by replacing its mean calculation with the mean calculated via low-rank factorization (as described in Equation 3). The upper confidence bound calculation and next example selection is the same as the original UCB-E algorithm.
The detailed algorithmic description can be found in Algorithm 6. We see that on the hard dataset GSM8K Prompts where UCB-E-LRF has certain benefit over UCB-E, there is a significant gap between UCB-E-LRF and UCB-E-LRF (Score Only). This means that the component of uncertainty matrix $R$ in UCB-E-LRF is crucial to contribute the benefit over UBC-E.
@@
Intuitively, a smaller batch size is more flexible can have more fine grained selection. We experiment with $b\in\{1,32,128\}$ and as shown in the plot, smallest batch size gives the best performance especially on harder datasets whereas a batch size of 128 significantly degrades the performance. However, it incurs more overall computational cost due to fitting low rank factorization more often than a smaller batch size; for example when $b=2$, it takes almost 16 times as much time as $b=32$ to achieve $1.0$ precision, which is ineffective.
Ensemble Size $S$ Ablation.
We vary the ensemble size $S\in\{4,16,64,256\}$ and find that very small ensemble size gives suboptimal performance on hard datasets. There is almost no performance difference on easy dataset like AlpacaEval (Drop Annotator). The performance is robust to different $S$ as long as it is larger than $64$.
@@
We experiment with different uncertainty scaling values $\{0.1,0.5,1,5,10\}$. In Figure 4d, it can be seen that a certain range of $\eta$, from $0.1$ to $5$, gives similar performance on both two datasets. This demonstrate the robustness of selecting $\eta$.
@@
Our algorithm UCB-E-LRF by default randomly evaluates $5\%$ of the method-example pairs in the data matrix before actively selects. We analyze the effect of varying the budget $T_{0}$ among $\{0.5\%,1\%,5\%,10\%\}\times n\times m$ on the algorithm performance on the two datasets. We see that a very small warm-up budget with $0.5\%$ of data can achieve decent precision initially, but fall behind compared to larger warm-up budget as more data are evaluated. In contrast, a very large warm-up budget of $10\%$ delays the active selection algorithm too much and also achieves suboptimal performance. We therefore use $5\%$ as a generally strong starting point. Note that on AlpacaEval, it is possible to achieve the same performance with even smaller warm-up budget, suggesting more saving is possible.
@@
Hyperparameter $r$ adjusts the bias-variance trade-off. Empirically we experiment with $r\in\{1,2,5\}$ and find that $r=1$ is consistently better than larger value counterparts. The results can be explained by the fact that a larger $r$ requires more evaluated data in order to prevent overfitting, which might not have advantage in the limited budget setting. Therefore, we recommend using $r=1$ for all datasets.
@@ 5 Related Work
The goal of best-arm identification [6, 2] is to find the arm with the highest reward by pulling these arms and getting the feedback. By making an analogy, in our problem method $f_{i}$ is the arm and the score $\mu_{i}$ of $f_{i}$ is the reward. There are two ways to define the best-arm identification problem: fixed budget and fixed confidence. In the fixed budget setting, the budget for the arm pulls is fixed and the algorithm is designed for better chances to identify the correct best arm – our problem defined in this paper has the similar evaluation budget. UCB-E [2] and Successive Elimination [2] are two pioneering algorithms proposed for this setting, followed by a line of improvement [30, 38, 10, 39]. Qin [51] states the optimal fixed budget best-arm identification as an open problem. In the setting of fixed confidence, the algorithms [35, 37, 31] work towards fewer number of arms to guarantee the given confidence of getting the best arm. Garivier and Kaufmann [24] gives an optimal algorithm in terms of the minimum of arms to pull. Another extension beyond the setting of fixed budget or confidence is the PAC learning framework, where the target is to maximize the chance of getting an mostly best arm, with a tolerance of $\varepsilon$ gap to the highest reward [36, 31, 12].
@@
Concurrently, Shi et al. [56] explores using multi-arm bandit for prompt optimization in a limited budget setting. Our work does not solely focus on prompt optimization and instead a more general evaluation setup. We also propose an algorithm that leverages low-rank factorization for better best-arm identification performance.
@@
As shown in the objective function of Equation 1, low-rank factorization is a non-convex optimization problem. Therefore a line of work focus on how to solve this non-convex optimization [9, 8, 14], while another line of work [13, 7] study the approximation error between the estimated low-rank matrix and the target matrix in terms of $p$ when assuming the observations are i.i.d. sampled with a pre-assumed chance $p$ and the additive noise to each observation is i.i.d. Gaussian noise.
@@
Tremendous effort has been devoted to developing effective evaluation functions to assess the quality of open-ended generations from language models. Early work in this direction such as BLEU [50] and ROUGE [46] are rule-based that use lexical overlaps to quantify similarity between a generated response and reference. However, lexical overlaps may not align well with the underlying semantics of the text. The shortcomings of these rule-based evaluation functions motivated a line of work studying using language models [65, 55, 63] to evaluate generations. A seminal work in this area is BERTScore which uses embeddings from a BERT model [21] to compute similarity. More recently, LLM-as-a-Judge [66] proposes using instruction-tuned large language models to evaluate generations. Zheng et al. [66] finds that costly proprietary models such as GPT4 have high agreement rate with human.
@@
Diverse benchmarks have been developed to evaluate LLM performance across various domains including natural language understanding on translations and sentiment analysis [4, 54, 57], mathematical and common sense reasoning [18, 29, 3], text retrieval and question answering [61, 33, 20]. We refer readers to [11] for a more comprehensive discussion on existing benchmarks. The commonality of these benchmarks is that a ground truth answer is typically given and once the responses from LLMs are generated, they can be evaluated fairly efficiently without incurring large amount of overhead such as money and compute. Recently, a new type of benchmark or more precisely leaderboard has been created attempting to compare various LLMs to each other. The typical setup involves providing the same prompt to two LLMs and then having their responses compared side by side with another LLM, which determines which response is considered superior and thus wins. Two notable leaderboards are AlpacaEval [44] and Chatbot Arena [15]. AlpacaEval has a static prompt dataset and for each prompt and LLM, the response is compared with that of a baseline model (GPT3 or GPT4). For Chatbot Arena, the prompts are submitted by users and two random LLMs are selected to respond to the user-written prompt. For both benchmarks, to aggregate the overall performance of LLMs, average win rate or ELO score can be calculated from the pairwise comparison statistics. Since the evaluation requires using another LLM to decide the better response, the evaluation incurs additional compute and / or monetary cost than the benchmarks that evaluate models based on ground truth answers.
6 Discussion and Conclusion
One limitation of our proposed algorithms is that we assume there is a two-dimensional fixed-size matrix as our dataset for active method-example selection. In some real-world applications, new rows or columns can be gradually incorporated and the matrix size is dynamic. In other scenarios such as Chatbot Arena, instead of being able to decide what examples to select, we can only select a pair of models to compare for a user-specified example. We leave studying how to minimize budget to find the best method for these new settings as future work.
In this work we focus on identifying the top-performer with limited budget as it is a very common problem; covering a wide range of applications particularly for practitioners who simply want to find the best prompt or model for deployment purposes. Another interesting direction for future work is recovering a full or top K ranking of performers, as it provides more fine-grained information that is necessary for specific use cases such as iterating on LLM design, and tracking LLM ranking over time. Of course in this setting the opportunity for resource savings is less than in the top-performer setting. In Figure 3, we evaluate our algorithms on NDCG@K as preliminary evidence supporting the possibility of resource saving recovering a top K ranking. We leave a more thorough study for future work.
In conclusion, we formulate the problem that when the evaluation for each model-example is resource intensive in regard to money, compute and time, given a certain budget of resource, how can we still have a high probability of identifying the best model on a specific dataset among a set of candidates. We propose two algorithms to tackle the problem, which all sequentially decide which model-example pair to evaluate next by observing the previous evaluated pairs. The first algorithm follows the idea of a classic multi-arm bandit algorithm UCB-E and enjoys the similar theoretical guarantee to lower bound the chance given any budget. The second algorithm, UCB-E-LRF extends UCB-E by leveraging the finding of approximate low-rank of the target evaluation matrix. In the experiment, we show that both algorithms introduced in our paper are significantly better than one that just uniformly samples method-example pairs. Moreover, we identified the condition when the UCB-E or UCB-E-LRF works better than the other. We do not foresee any negative societal impacts for our work.
7 Acknowledgement
We thank Justin Lovelace, Varsha Kishore, and Katie Luo for their helpful discussion and feedback on the paper draft. JPZ is supported by a grant from the Natural Sciences and Engineering Research Council of Canada (NSERC) (567916). CKB and CPG are supported by the Cornell University AI for Science Institute, the AI Climate Institute, the National Science Foundation, the National Institute of Food and Agriculture, and the Air Force Office of Scientific Research. RW is supported by grants from National Science Foundation NSF (CIF-2402817, CNS-1804829), SaTC-2241100, CCF-2217058, ARO-MURI (W911NF2110317), and ONR under N00014-24-1-2304. Wen Sun is supported by NSF IIS-2154711 and NSF CAREER 2339395. This research is also supported by grants from the National Science Foundation NSF (IIS-2107161, and IIS-1724282, HDR-2118310), the Cornell Center for Materials Research with funding from the NSF MRSEC program (DMR-1719875), DARPA, arXiv, LinkedIn, and the New York Presbyterian Hospital.
Appendix A Additional Details - Datasets

### v3

@@ 4.4 Main Results
In Figure 3, we plot the performance of baselines and our algorithms on all six datasets (columns) as the budget $T$ increases from 5% to 100% of the total number of method-example pairs. In each row, we evaluate the algorithms on a different metric (either top 1 precision with different $\epsilon$, $p$ or NDCG@10). Each line in the figure is the average result over 50 independent trials with different seeds. For example, an average top 1 precision of 0.9 indicates that 45 out of 50 trials predict the best method correctly at the budget level that its x-coordinate represents. In addition, we calculate $H_{1}$ as defined in Corollary 1 that quantifies the difficulty of finding the best method on a dataset. Intuitively, a higher $H_{1}$ value suggests that the method set $\mathcal{F}$ is large or there are many methods that have similar performance with the best one and distinguishing them can be challenging. In Appendix C Figure 5, we also plot the histogram to show the distribution of performance among $\mathcal{F}$ on these datasets.
@@
As seen from Figure 3, both UCB-E and UCB-E-LRF consistently achieve high precision and NDCG with much less budget compared to the baselines. For example, on AlpacaEval (Drop Annotator), our proposed algorithms can reach a precision of 1 with just 8% budget whereas the baselines require 80-90%, an order of magnitude more budget needed. These results suggest that it is entirely possible to identify the best method without exhaustively evaluating all method-example pairs. They further demonstrate that the active selection algorithms (our two algorithms) are more efficient than the non-active algorithms (three baselines). Additionally, the better NDCG performance from our proposed algorithms shows that our methods can more correctly rank top-performance methods.
@@ 4.5 More Empirical Analysis
Yes, going through Figure 3 from left to right, as $H_{1}$ decreases, the percentage of matrix evaluation needed to reach a precision of 1 also decreases from more than 20% to just under 5%. Moreover, the $H_{1}$ values seem to be related to the tasks. Prompt engineering datasets typically have higher $H_{1}$ possibly due to the homogeneity of prompt performance with the same LLM. In contrast, datasets that benchmark different LLMs such as AlpacaEval make it much easier to find the best performing model.
@@
To study the effect of actively selecting both the next method and next example to evaluate using uncertainty matrix $R$, we consider an ablation of UCB-E-LRF. In this variant, called UCB-E-LRF (Score Only), we modify the UCB-E algorithm by replacing its mean calculation with the mean calculated via low-rank factorization (as described in Equation 3). The upper confidence bound calculation and next example selection are the same as the original UCB-E algorithm.
The detailed algorithmic description can be found in Algorithm 6. We see that on the hard dataset GSM8K Prompts where UCB-E-LRF has a certain benefit over UCB-E, there is a significant gap between UCB-E-LRF and UCB-E-LRF (Score Only). This means that the component of uncertainty matrix $R$ in UCB-E-LRF is crucial to contributing the benefit over UBC-E.
@@
Intuitively, a smaller batch size is more flexible and can have more fine grained selection. We experiment with $b\in\{1,32,128\}$, and as shown in the plot, the smallest batch size gives the best performance, especially on harder datasets whereas a batch size of 128 significantly degrades the performance. However, it incurs more overall computational cost due to fitting low rank factorization more often than a smaller batch size; for example when $b=2$, it takes almost 16 times as much time as $b=32$ to achieve $1.0$ precision, which is ineffective.
Ensemble Size $C$ Ablation.
We vary the ensemble size $C\in\{4,16,64,256\}$ and find that very small ensemble size gives a suboptimal performance on hard datasets. There is almost no performance difference on easy datasets like AlpacaEval (Drop Annotator). The performance is robust to different $C$ as long as it is larger than $64$.
@@
We experiment with different uncertainty scaling values $\{0.1,0.5,1,5,10\}$. In Figure 4d, it can be seen that a certain range of $\eta$, from $0.1$ to $5$, gives similar performance on both two datasets. This demonstrates the robustness of selecting $\eta$.
@@
Our algorithm UCB-E-LRF by default randomly evaluates $5\%$ of the method-example pairs in the data matrix before selecting actively. We analyze the effect of varying the budget $T_{0}$ among $\{0.5,1,5,10\}\%$ on the algorithm performance on the two datasets. A very small warm-up budget with $0.5\%$ of data can achieve decent precision initially, but fall behind compared to larger warm-up budget as more data are evaluated. In contrast, a very large warm-up budget of $10\%$ delays the active selection algorithm too much and also achieves suboptimal performance. We therefore use $5\%$ as a generally strong starting point. On AlpacaEval, it is possible to achieve the same performance with an even smaller warm-up budget, suggesting more savings is possible.
@@
Hyperparameter $r$ adjusts the bias-variance trade-off. Empirically we experiment with $r\in\{1,2,5\}$ and find that $r=1$ is consistently better than larger value counterparts. The results can be explained by the fact that a larger $r$ requires more evaluated data in order to prevent overfitting, which might not have an advantage in the limited budget setting. Therefore, we recommend using $r=1$ for all datasets.
@@ 5 Related Work
The goal of best-arm identification (Bubeck et al., 2009; Audibert and Bubeck, 2010) is to find the arm with the highest reward by pulling these arms and getting feedback. By making an analogy, in our problem method $f_{i}$ is the arm and the score $\mu_{i}$ of $f_{i}$ is the reward. There are two ways to define the best-arm identification problem: fixed budget and fixed confidence. In the fixed budget setting, the budget for the arm pulls is fixed and the algorithm is designed for better chances to identify the correct best arm – our problem defined in this paper has a similar evaluation budget. UCB-E (Audibert and Bubeck, 2010) and Successive Elimination (Audibert and Bubeck, 2010) are two pioneering algorithms proposed for this setting, followed by a line of improvement (Honda and Takemura, 2011; Kaufmann et al., 2012; Cappé et al., 2013; Kaufmann et al., 2016). Qin (2022) states the optimal fixed budget best-arm identification as an open problem. In the setting of fixed confidence, the algorithms (Kalyanakrishnan et al., 2012; Kaufmann and Kalyanakrishnan, 2013; Jamieson et al., 2014) work towards fewer number of arms to guarantee the given confidence of getting the best arm. Garivier and Kaufmann (2016) gives an optimal algorithm in terms of the minimum of arms to pull. Another extension beyond the setting of fixed budget or confidence is the PAC learning framework, where the target is to maximize the chance of getting an mostly best arm, with a tolerance of $\varepsilon$ gap to the highest reward (Karnin et al., 2013; Jamieson et al., 2014; Chaudhuri and Kalyanakrishnan, 2019).
@@
Concurrently, Shi et al. (2024) explore using multi-arm bandit for prompt optimization in a limited budget setting. Although there are similarities in the approach, it is fair to say that their work differs substantially in focus. In contrast, our main goal is to speed up model evaluation rather than prompt optimization. Further, we also propose an algorithm that leverages low-rank factorization for better best-arm identification performance. We discuss additional related work on LLM performance evaluation functions and benchmarks in Appendix A, which are relevant but orthogonal to our problem.
6 Discussion and Conclusion
In this work we focus on identifying the top-performer with a limited budget as it is a very common problem; covering a wide range of applications particularly for practitioners who simply want to find the best prompt or model for deployment purposes. One interesting direction for future work is recovering a full or top K ranking of performers, as it provides more fine-grained information that is necessary for specific use cases such as iterating on LLM design, and tracking LLM ranking over time. Of course the opportunity for resource savings is less than the top-performer setting. In Figure 3, we evaluate our algorithms on NDCG@K as preliminary evidence supporting the possibility of resource saving recovering a top K ranking. We leave a more thorough study as future work.
A limitation of our proposed algorithms is that we assume there is a two-dimensional fixed-size scoring matrix for active method-example selection. In some real-world applications, new rows or columns can be gradually incorporated, making the matrix size dynamic. In other scenarios such as Chatbot Arena, instead of being able to decide what examples to select, we can only select a pair of models to compare for a user-specified example. We leave studying how to efficiently select the best method in these settings as another direction for future work.
In conclusion, we tackle the challenge of resource-constrained evaluation, where assessing method-example pairs is costly in terms of money, compute, and time. Our goal is to identify the best method with high probability while staying within budget. We propose two sequential decision-making algorithms: UCB-E, inspired by the classic multi-arm bandit approach, which provides a theoretical guarantee on success probability, and UCB-E-LRF, which extends UCB-E by leveraging low-rank factorization to approximate the scoring matrix. Experiments show both outperform random sampling, and we identify conditions where UCB-E or UCB-E-LRF excels.
7 Reproducibility Statement
Our code is available at https://github.com/kilian-group/banditeval.
Acknowledgement
We thank Justin Lovelace, Varsha Kishore, and Katie Luo for their helpful discussion and feedback on the paper draft. JPZ is supported by a grant from the Natural Sciences and Engineering Research Council of Canada (NSERC) (567916). CKB and CPG are supported by Schmidt Sciences programs, an AI2050 Senior Fellowship and two Eric and Wendy Schmidt AI in Science Postdoctoral Fellowships; the National Science Foundation (NSF) including the NSF Research Traineeship program; the National Institute of Food and Agriculture; and the Air Force Office of Scientific Research. RW is supported by grants from NSF (CIF-2402817, CNS-1804829), SaTC-2241100, CCF-2217058, ARO-MURI (W911NF2110317), and ONR under N00014-24-1-2304. WS is supported by NSF IIS-2154711, NSF CAREER 2339395 and DARPA LANCER: LeArning Network CybERagents. This research is also supported by grants from the NSF (IIS-2107161, and IIS-1724282, HDR-2118310), the Cornell Center for Materials Research with funding from the NSF MRSEC program (DMR-1719875), DARPA, arXiv, LinkedIn, and the New York Presbyterian Hospital.
Appendix A Further Discussion on Related Work
@@
As shown in the objective function of Equation 1, low-rank factorization is a non-convex optimization problem. Therefore a line of work focus on how to solve this non-convex optimization [Candès and Tao, 2010, Candes and Recht, 2012, Chi et al., 2019], while another line of work [Chen et al., 2020, Cai et al., 2019] study the approximation error between the estimated low-rank matrix and the target matrix in terms of $p$ when assuming the observations are i.i.d. sampled with a pre-assumed chance $p$ and the additive noise to each observation is i.i.d. Gaussian noise.
@@
Tremendous effort has been devoted to developing effective evaluation functions to assess the quality of open-ended generations from language models. Early work in this direction such as BLEU [Papineni et al., 2002] and ROUGE [Lin, 2004] are rule-based that use lexical overlaps to quantify similarity between a generated response and reference. However, lexical overlaps may not align well with the underlying semantics of the text. The shortcomings of these rule-based evaluation functions motivated a line of work studying using language models [Zhang et al., 2019, Sellam et al., 2020, Yuan et al., 2021] to evaluate generations. A seminal work in this area is BERTScore which uses embeddings from a BERT model [Devlin et al., 2018] to compute similarity. More recently, LLM-as-a-Judge [Zheng et al., 2024] proposes using instruction-tuned large language models to evaluate generations. Zheng et al. [2024] finds that costly proprietary models such as GPT4 have high agreement rate with human.
@@
Diverse benchmarks have been developed to evaluate LLM performance across various domains including natural language understanding on translations and sentiment analysis [Bojar et al., 2016, Sahayak et al., 2015, Singh et al., 2024], mathematical and common sense reasoning [Cobbe et al., 2021, Hendrycks et al., 2021c, Bisk et al., 2020], text retrieval and question answering [Yang et al., 2015, Jin et al., 2019, Dell et al., 2023]. We refer readers to Chang et al. [2024] for a more comprehensive discussion on existing benchmarks. The commonality of these benchmarks is that a ground truth answer is typically given and once the responses from LLMs are generated, they can be evaluated fairly efficiently without incurring large amount of overhead such as money and compute. Recently, a new type of benchmark or more precisely leaderboard has been created attempting to compare various LLMs to each other. The typical setup involves providing the same prompt to two LLMs and then having their responses compared side by side with another LLM, which determines which response is considered superior and thus wins. Two notable leaderboards are AlpacaEval [Li et al., 2023] and Chatbot Arena [Chiang et al., 2024]. AlpacaEval has a static prompt dataset and for each prompt and LLM, the response is compared with that of a baseline model (GPT3 or GPT4). For Chatbot Arena, the prompts are submitted by users and two random LLMs are selected to respond to the user-written prompt. For both benchmarks, to aggregate the overall performance of LLMs, average win rate or ELO score can be calculated from the pairwise comparison statistics. Since the evaluation requires using another LLM to decide the better response, the evaluation incurs additional compute and / or monetary cost than the benchmarks that evaluate models based on ground truth answers.
Appendix B Derivation of Corollary 1
By mostly following the proof of the original theorem of UCB-E (Section 6.2 in Audibert and Bubeck [2010]), we only need to show that
$\mathbb{P}\left(\left|\sum_{j=1}^{n}\frac{O_{ij}\cdot S_{ij}^{\rm obs}}{\sum_{j=1}^{n}O_{ij}}-\frac{\sum_{j=1}^{n}S_{ij}}{n}\right|<\frac{1}{5}\sqrt{\frac{a}{\sum_{j=1}^{n}O_{ij}}}\right)\geq 1-\exp\left(-\frac{2a}{25}\right).$
This is equivalent to prove if there are $p$ random variables $\{X_{1},\cdots,X_{p}\}$ sampling without replacement from the finite set $\{S_{i1},\cdots,S_{in}\}$,
$\mathbb{P}\left(\left|\frac{X_{1}+\cdots+X_{p}}{p}-\frac{\sum_{j=1}^{n}S_{ij}}{n}\right|<\frac{1}{5}\sqrt{\frac{a}{p}}\right)\geq 1-\exp\left(-\frac{2a}{25}\right).$
This can be implied by the analogous Hoeffding inequality [Hoeffding, 1994] that works for the sum of random variables sampling without replacement.
Appendix C Additional Details - Datasets